OpenStudy (anonymous):

*in beats per minute * using a heart rate formula that is 0.6(220-x) and X representing the persons age, if a heart rate is 115 beats per minute what is the age in years (rounded to nearest year)

OpenStudy (anonymous):

R = 0.6(220-x) 115 = 0.6(220-x) 115/0.6 = 220 - x now, solve for x and round to nearest year. Answer is between 25 and 30.